OpenStudy (anonymous):

I need a quick help!!! If u have this sign > or equal to do u shade above or below the line? Please help!! Its an easy way to get a MEDAL :)))

OpenStudy (anonymous):

below

OpenStudy (anonymous):

\[\ge\] like this

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

So u go below? U sure???

OpenStudy (anonymous):

@matt.mawson sure?

OpenStudy (anonymous):

It is below

OpenStudy (anonymous):

He is right. You go below.

OpenStudy (anonymous):

\[\ge\]

OpenStudy (anonymous):

Alright thanx guys!!!!

OpenStudy (anonymous):

But if I do that in my graph it woukdnt make sense

OpenStudy (anonymous):

@dmot7291 & @courtney_lee

OpenStudy (anonymous):

Can you draw the graph, what do you mean?

OpenStudy (anonymous):

How do I draw it for u where do I go?

OpenStudy (anonymous):

Click draw below the reply box at the bottom

OpenStudy (anonymous):

I dont c nothing but the box to type

OpenStudy (anonymous):

Can you give me the equation you have, I think i understand your question. You are trying to plot an inequality right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok but dont panic

OpenStudy (anonymous):

If it is less than or equal too you go below. If it is greater than or equal too you go above.

OpenStudy (anonymous):

x >=0 y>=0 x + 2y =< 6 2y - x =< 2 x + y =< 5 f (x, y) = 3x - 5y That's the question its from an algebra 2 book

OpenStudy (anonymous):

Y>0 would be shaded above because y is greater than 0. y< is below because y is less than

OpenStudy (anonymous):

http://my.hrw.com/math06_07/nsmedia/tools/Graph_Calculator/graphCalc.html online graphing calculator, can show shading too

OpenStudy (anonymous):

Thanx :)

OpenStudy (anonymous):

Nah it didnt work

OpenStudy (anonymous):

What are you trying to plot exactly?

OpenStudy (anonymous):

3rd equation

OpenStudy (anonymous):

solve for y before you graph it

OpenStudy (anonymous):

Will it be this at the end y = -x + 3?

OpenStudy (anonymous):

Anybody?

OpenStudy (anonymous):

3rd equation is y =< -(1/2)x + 3

OpenStudy (anonymous):

Right forgot to put 1/2 thanx